Could someone please help me with this word problem? I can't get past the point of the inverse matrix.
A local sorority sold hotdogs and bratwursts at the spring fling picnic. The first day they sold 8 dozen hotdogs and 13 dozen bratwursts for 339.60. The second day they sold 10 dozen hotdogs and 15 dozen bratwursts for a total of 402.00. How much did each cost?
Each hotdog sold for $?
Each bratwurst sold for $?

Answer by checkley77(12569)

8*12H+13*12B=339.60
96H+156B=339.60 (MULTIPLY BY 5)
10*12H+15*12B=402.00
120H+180B=402.00 (MULTIPLY BY-4 & ADD)
480H+780B=1698
-480H-720B=-1608
----------------------------
0H+60B=90
B=90/60
B=$1.50 FOR EACH BRATWURST.
10*12H+15*12*1.5=402
120H+270=402
120H=402-270
120H=132
H=132/120
H=$1.10 FOR EACH HOTDOG.
PROOF:
10*12*1.10+15*12*1.5=402
120*1.10+180*1.5=402
132+270=402
402=402